OpenStudy (ashleyisakitty):

Find the missing value to the nearest hundredth. tan__=26 None of my answer choices align with 26? help?

OpenStudy (ashleyisakitty):

The answer choices are: 89.08, 87.80, 64.56 and 61.94 but according to my calculations none of them fit with 26

OpenStudy (freethinker):

maybe it is an inverse tangent = 26

OpenStudy (anonymous):

Make sure your calculator is in degree mode.

OpenStudy (freethinker):

and of course that too, like what @SACAPUNTAS said

OpenStudy (ashleyisakitty):

I dont have a graphing calculator, i only use wolffram.

OpenStudy (freethinker):

remember that tangent means the ratio of y and x y/x

OpenStudy (ashleyisakitty):

?? any further assistance? :x How can i figure it out without being able to turn my graping into degrees?

OpenStudy (shamil98):

tan 87.80 = 26

OpenStudy (ranga):

Put in 26 where it says arctan and click on = http://www.rapidtables.com/calc/math/Arctan_Calculator.htm

OpenStudy (shamil98):

I used my scientific calculator, and that's the result. @Ashleyisakitty

OpenStudy (shamil98):

wolfram has it too, just dont click enter ;P

OpenStudy (ashleyisakitty):

really? i got -0.1

OpenStudy (shamil98):

it gives the quicksearch answer i believe.

OpenStudy (shamil98):

if you dont press enter

OpenStudy (ashleyisakitty):

sometimes it does, not for this in particular, thanks everyone.

OpenStudy (ashleyisakitty):

Any ideas on what sin__=8/40 would be? How to get it?

OpenStudy (shamil98):

sin x = 8/40?

OpenStudy (ashleyisakitty):

yes

OpenStudy (ranga):

8/40 = 0.2. Look up inverse sign of 0.2

OpenStudy (shamil98):

yeah

OpenStudy (ranga):

i meant sin not sign

OpenStudy (shamil98):

sin ^-1 = 8/40 sin = 11.5369

OpenStudy (ranga):

answer in degrees

OpenStudy (freethinker):

http://www.wolframalpha.com/input/?i=arctan+26+degrees

OpenStudy (freethinker):

like I said earlier, inverse tangent (arctan)